# On the approximability of random-hypergraph MAX-3-XORSAT problems with quantum algorithms

### v1

@@ IV.4 Achievable approximation ratio with spectrally folded trial minimum annealing
With this result in hand, we will now predict the macroscopic quantum tunneling rate–and thus, achievable approximation ratio–for the spectrally folded Hamiltonian. From this, we can calculate our target value of $A$, assuming that there is a single deep minimum far below the energy of any local minima. Relaxing this assumption will improve the performance of the algorithm by virtue of there being many more target states, as confirmed in our simulations for lower constraint densities. We consider the protocol in section III, with an initial state $\left|L\right>$. We assume that the process of ramping the transverse field up and down is itself at least roughly adiabatic, i.e., we can assume approximate spectral continuity with respect to the folding and lowering Hamiltonians, noting that the lowering Hamiltonian will itself create $O\left(\sqrt{N}\right)$ shifts to the energies of states near the fold. It follows from our assumption that the ramping process itself does not meaningfully heat the system. We then consider the set $\mathcal{T}$ of all states within $O\left(1\right)$ shifts of $-AE_{\mathrm{GS}}$ in $H_{P}$, the states closest to the fold, and compute, as a function of all our various algorithm parameters, the total probability of tunneling into any one of them.
Since the tunneling rate into any individual state is exponentially small, and the time over which we slowly turn off the lowering Hamiltonian is $T\propto O\left(N\right)$, we can assume that tunneling will be diabatic with respect to any individual state. A Fermi’s Golden rule analysis as in Kapit and Oganesyan (2021) suggests that
$P_{\mathrm{tot}}\propto\frac{T}{W}\sum_{j\in\mathcal{T}}\Omega_{0,Lj}^{2},$ (29)
@@
where $W\sim O\left(N\right)$ is the energy range swept over by reducing $C\left(t\right)$ to 0, and $\Omega_{0,Lj}$ is the tunnel splitting at degeneracy between $\left|L_{D}\right>$ and the target state $\left|j_{D}\right>$, which we assume are an average of $\sim N/2$ flips apart. To go further, we need to compute the average value of $\Omega_{0,Lj}^{2}$, noting that while of course there will be substantial state-to-state variations, given that there are exponentially many states in $\mathcal{T}$ the average value is going to dominate Eq. (29). As in the previous calculation, the most important quantity here is the average cost per flip away from the typical ground state in the folded Hamiltonian, which remarkably turns out to be $A$-independent.
To see this, we start from the average cost per flip away from $\left|G\right>$, given by Eq. (11), and note that we can invert that equation to find the mean number of flips $x_{A}N$ for which $E\left(x_{A}N\right)\right>=-AN$. To be specific,
$x_{A}=\frac{1-A^{1/3}}{2}.$ (30)
We can therefore assume that the typical state in $\mathcal{T}$ is $x_{A}N$ flips away from $\left|G\right>$. If we consider the sequences of primary spin flips connecting $\left|j_{D}\right>$ and $\left|L_{D}\right>$, the typical flip sequence starts $x_{A}N$ flips away from $\left|G\right>$, and notice that with probability $1-x_{A}$ an additional random flip towards $\left|L\right>$ will also move closer to $\left|G\right>$. Taking all these effects and the division by $A$ into account, so that the bare unperturbed energy of both $\left|j\right>$ and $\left|L\right>$ when they cross are both $\sim-N$, a bit of algebra shows that for $y$ flips away from a ground state of the folded Hamiltonian, not only is the total average cost $\Delta E\left(y\right)$ $A$-independent, it is precisely equal to the cost given by Eq. (11).
@@
This is again only an average, but noting that since it appears the denominators of equations like (IV.3), variations about it are more likely to increase the tunneling rate than decrease it. And likewise, since $H_{L}$ is a random 3-XORSAT problem itself the mean cost per flip away from $\left|L\right>$ is going to be given by Eq. 11 as well, so Eqns. IV.3 through IV.3 can faithfully predict the average tunneling rate between $\left|L\right>$ and a randomly chosen ground state of the folded Hamiltonian. [^5] We expect that using the average cost per flip in Eq. 17 will if anything underestimate the per-state tunneling rate in real disordered problems. This is because all of these energy costs appear in denominators, which leads to likely small asymmetries in how much the deviations from the average in any individual flip sequence contribute to the total matrix element, giving lower energy sequences proportionally higher weight. We do not really expect this effect to be significant but rather highlight it as another point where our prediction is conservative by design. For more discussion of this approximation, see the appendix.
@@
The achievable approximation ratio is thus determined by the per-state decay exponent $b$, computed in section IV.3 as a function of $N$ and $\kappa$ by fitting Eq. (IV.3) to $\Omega_{0}\left(N\right)\propto\sqrt{N}2^{-bN}$, and then choosing $x_{A}$ using Eq. (30) to solve Eq. (IV.4). This analysis only counts states within $O\left(1\right)$ shifts of $AE_{\mathrm{GS}}$ (recall that $E_{\mathrm{GS}}=-N$ in our normalization) and ignores low-order polynomial prefactors; for $b=0.2$, which again depends on $\kappa=1.29$ in this calculation, this is solved when $x_{A}\simeq 0.08$, or $A\simeq 0.59$.
@@
This means that if the true ground state satisfies a constraint fraction $F$ beyond random guessing—e.g. a total fraction $1/2+F$, so $F$ is at most 1/2 here—our algorithm will return states which satisfy a fraction $1/2+AF$ with high probability. If we choose $A$ to be too large compared to the target value established by expressions like Eq. (IV.4), we risk failing to well-approximate the problem; conversely, choosing $A$ below it will reduce the returned approximation ratio to $A$ and thus perform suboptimally. And, we emphasize again, this prediction assumes a random, potentially dense hypergraph but is fundamentally independent of the fraction satisfied in $E_{\mathrm{GS}}$ itself and so applies to the planted partial solution instances we use for numerical benchmarking below.
IV.5 Further Comments and Caveats
@@
We expect that this analysis underestimates the choice of $A$ that will return states with $E\leq AE_{\mathrm{GS}}$ with constant probability. This is because our counting here only counts states very close to the fold, when in reality the probability of tunneling into states a small extensive fraction larger than $AE_{\mathrm{GS}}$ is still going to be appreciable due to the continued exponential growth of the number of targets, even if the per-state tunneling rate does tend to decrease with increasing $E$ due to the interference effect mentioned earlier, in which perturbative corrections that mix with states of lower energy have opposite sign. In addition to this consideration, because the target ground states and low lying excitations of the folded Hamiltonian in $\mathcal{T}$ very roughly form a hyperspherical shell, any individual target state will have other states in $\mathcal{T}$ that are relatively close to it in Hamming distance. Consequently we expect these states to have a band dispersion, centered around the mean energy given by the corrections in Eq. (IV.3). Since we are already assuming off-resonant tunneling, i.e. a per-state tunneling rate $\propto\Omega_{0}^{2}$, and so summing over squared matrix elements, if we consider states near the band center where the density is highest this will alter the average tunneling rate by at most a prefactor. However, there are good reasons to suspect that tunneling into extremal states near the bottom of the band can be substantially enhanced, enough to increase the optimal value of $A$. This calculation is difficult to do quantitatively so we do not attempt it here; we instead see the approximation of considering only states near the band center of $\mathcal{T}$ as another choice that likely underestimates the achievable approximation ratio. We discuss this point in more detail in Appendix B.
@@
For quadratic folding AQC as in Eq. (6), if we choose $A=1$ the gap is efficiently computable using the methods in Jörg et al. (2010) and decays as $\Omega_{0}\sim 2^{-0.16N}$. Given that, like linear folding, the cost per flip curve is $A$-independent, if we assume that the tunneling rate per state for $A<1$ is basically equal to this, then the total decay exponent vanishes if $x_{A}\simeq 0.06$ and $A\simeq 0.68$ using the arguments of the previous few paragraphs. We do not think that can be simply assumed as easily as with tunneling between semiclassical minima and a linearly folded problem Hamiltonian, in the DPP, and more theoretical work is needed here to analytically determine the optimal choice of $A$. Interestingly however, our simulation data in Sec. VI supports this conclusion, with a worst case polynomial time approximation ratio of 0.7 found in our simulations. These simulations show that this method performs similarly too, or slightly worse than, the 3XORSAT-TMA algorithm, which is better able to outperform the approximation guarantee of $\sim 0.6$ derived here. [^6] For smaller systems we also tested linear folding AQC and quadratic folding TMA. In very preliminary studies we found that quadratic AQC modestly outperformed linear AQC, and linear TMA more significantly outperformed quadratic TMA. So we chose not to pursue those methods for larger simulations and do not present those results here, but they may be viable or even superior for other problem classes. We also expect that the average tunneling rate–and thus, achievable approximation ratio–can likely be further increased by using other, potentially many-frequency, AC methods such as RFQA Kapit (2021b); Kapit and Oganesyan (2021); Tang and Kapit (2021); Mossi et al. (2023); Grattan et al. (2023). For simplicity, we do not incorporate these methods in this work, but they could be a novel way to further improve the performance of this algorithm and are worth exploring in future research.
@@
In summary, through a relatively novel resummed extensive order perturbation theory, we have shown that random hypergraph MAX-3-XORSAT instances, including extremal ones with planted partial solutions, are efficiently approximable to a fairly large constant fraction through the spectrally folded quantum optimization algorithms. We do not expect this to be the case for QAOA, an expectation supported by the numerical evidence we present below. We similarly do not expect such guarantees to be possible for directly finding global optima, for the reasons set forth in the introduction. Evidently, one cannot so easily summit a mountain in hyperspace, but one can reach the rim of a crater. We now present a series of numerical simulations to further support these claims.
@@ V.1 Setup and summary of results
To confirm our predictions–or at least, verify that any serious issues with our calculations and interpretation of the problem are subtle and not apparent at system sizes within reach of present or near-future classical simulations–we performed a series of numerical simulations of various classical and quantum algorithms applied to our PPSPs. For all quantum simulation tasks we used the Qulacs package Suzuki et al. (2021). For smaller systems and algorithm prototyping we ran our simulations on local workstations; this includes all the spectral folding TMA simulations. For all QAOA and spectral folding AQC simulations presented, we used the Fujitsu Quantum Simulator, a classical HPC system. This allowed us to probe larger system sizes while still averaging over enough instances to have reliable statistics for the problem class. Unless otherwise stated, each datapoint represents the average over 960 or 1000 (for the QAOA and spectral folding AQC results) random problem instances constructed with the prescriptions outlined in section II.4. In all cases in this work we used an unsatisfied fraction $\epsilon=0.1$ for our partial planted solutions; we expect similar phenomenological behavior for other constant $\epsilon$ values, though we expect the precise thresholds we measure will vary with $\epsilon$ for classical optimization algorithms and QAOA (but not folded optimization, below its predicted performance floor).
@@
To determined the polynomial time hardness threshold, we applied a very simple rule where we fitted $P_{q}\left(N\right)$ to a simple exponential function, assuming any observed decay faster than $2^{-0.005N}$ corresponded asymptotically to exponential decay, and any positive exponents, e.g. exponential growth, represent small-$N$ growth toward some constant saturation value. This threshold of $2^{-0.005N}$ was chosen to reduce the influence of uncertainty from fitting in a small handful of cases. For the greedy algorithm with $N_{C}/N=1.5\sqrt{N}$ we fitted decay to an exponential in $\sqrt{N}$ as discussed below. The polynomial time approximation hardness threshold for a given algorithm, unsatisfied fraction $\epsilon$ and $N_{C}/N$ scaling choice is defined to be $q_{a}$, the largest values of $q$ for which we do not observe exponential decay. For $N_{C}/N=3\sqrt{N}/2$ this threshold is decaying with system size, and the asterisk next to the result for QAOA is to highlight the fact that we only ran these simulations out to $N=28$ so are likely not capturing the asymptotic threshold. As discussed elsewhere, the number of cost function calls is $O\left(N\right)$ for all approaches studied. Cases for which a range is quoted are where we felt there was some ambiguity to the fitting, and all values are the result of extrapolating fits to numerical simulations and are naturally somewhat approximate.
$N_{C}/N$ Classical QAOA AQC/0.75 AQC/0.85 TMA-3/0.75 TMA-3/0.85 TMA-L/0.75
2 0.75-0.8 0.75 0.75 0.75 0.75 0.75 0.7
4 0.55 0.55 0.7 0.7 0.75 0.8 0.75
6 0.45 0.45 0.75 0.7 0.8 0.8 0.75
3 $\sqrt{N}/2$ decaying 0.25/decaying* 0.75 0.75 0.8 0.8 0.75-0.8
@@
Table 1: Approximation hardness thresholds for the classical greedy search, high-depth QAOA and spectral folding variations. This table lists $q_{a}$, the largest value of the approximation ratio $q$ before exponential decay is reported, drawn from the numerical experiments in figures 5 through 8 and 14. Spectral folding results are labeled as protocol/$A$ (where $A$ is the approximation target); AQC is quadratic spectral folding in the AQC formulation, TMA-3 is trial minimum annealing with a linear folded Hamiltonian and 3-XORSAT lowering Hamiltonian, and TMA-L is the same with local $Z$ biases for the lowering Hamiltonian. These results support the predictions in section IV that random hypergraph problems are efficiently approximable through spectrally folded quantum optimization.
V.2 Performance of greedy classical algorithms

### v3

@@ IV.3.2 Achievable approximation ratio with spectrally folded trial minimum annealing
With this result in hand, we will now predict the macroscopic quantum tunneling rate–and thus, achievable approximation ratio–for the spectrally folded Hamiltonian. From this, we can calculate our target value of $A$, assuming that there is a single deep minimum far below the energy of any local minima. Relaxing this assumption should improve the performance of the algorithm by virtue of there being many more target states. We consider the protocol in section III.4, with an initial state $\left|L\right>$. We assume that the process of ramping the transverse field up and down is itself at least roughly adiabatic, i.e., we can assume approximate spectral continuity with respect to the folding and lowering Hamiltonians, noting that the lowering Hamiltonian will itself create $O\left(\sqrt{N}\right)$ shifts to the energies of states near the fold. It follows from our assumption that the ramping process itself does not meaningfully heat the system. We then consider the set $\mathcal{T}$ of all states within $O\left(1\right)$ shifts of $-AE_{\mathrm{GS}}$ in $H_{P}$, the states closest to the fold, and compute, as a function of all our various algorithm parameters, the total probability of tunneling into any one of them.
@@
Since the tunneling rate into any individual state is exponentially small, and the time over which we slowly turn off the lowering Hamiltonian is $T\propto O\left(N\right)$, we can assume that tunneling will be diabatic with respect to any individual state. A Fermi’s Golden rule analysis as in Kapit and Oganesyan (2021) suggests that the total success probability will be given by Eq. 2, where again $W\sim O\left(N\right)$ is the energy range swept over by reducing $C\left(t\right)$ to 0, and $\Omega_{0,Lj}$ is the tunnel splitting at degeneracy between $\left|L_{D}\right>$ and the target state $\left|j_{D}\right>$, which we assume are an average of $\sim N/2$ flips apart. To go further, we need to compute the average value of $\Omega_{0,Lj}^{2}$, noting that while of course there will be substantial state-to-state variations, given that there are exponentially many states in $\mathcal{T}$ the average value calculated in Eq. IV.3.1 is going to dominate Eq. (2).
@@
This is again only an average, but noting that the classical cost per flip for any given sequence appears the denominators of equations like (IV.3.1), variations about it are more likely to increase the tunneling rate than decrease it. And as remarked earlier, path-to-path variations about the average are suppressed by $O\left(1/\sqrt{d_{C}}\right)$ and thus naturally small in the approximation hard regime. And likewise, since $H_{L}$ is a random 3-XORSAT problem itself the mean cost per flip away from $\left|L\right>$ is going to be given by Eq. 18 as well, so Eqns. 27 through IV.3.1 can faithfully predict the average tunneling rate between $\left|L\right>$ and a randomly chosen ground state of the folded Hamiltonian. [^9] We expect that using the average cost per flip in Eq. 28 will if anything underestimate the per-state tunneling rate in real disordered problems. This is because all of these energy costs appear in denominators, which leads to likely small asymmetries in how much the deviations from the average in any individual flip sequence contribute to the total matrix element, giving lower energy sequences proportionally higher weight. We do not really expect this effect to be significant but rather highlight it as another point where our prediction is conservative by design. For more discussion of this approximation, see the appendix.
@@
The achievable approximation ratio is thus determined by the per-state decay exponent $b$, computed in section IV.3.1 as a function of $N$ and $\kappa$ by fitting Eq. (IV.3.1) to $\Omega_{0}\left(N\right)\propto\sqrt{N}2^{-bN}$, and then choosing $x_{A}$ using Eq. (27) to solve Eq. (IV.3.2). This analysis only counts states within $O\left(1\right)$ shifts of $AE_{\mathrm{GS}}$ (recall that $E_{\mathrm{GS}}=-N$ in our normalization) and ignores low-order polynomial prefactors; for $b=0.2$, which again depends on $\kappa=1.29$ in this calculation, this is solved when $x_{A}\simeq 0.08$, or $A\simeq 0.59$.
@@
This means that if the true ground state satisfies a constraint fraction $F$ beyond random guessing—e.g. a total fraction $1/2+F$, so $F$ is at most 1/2 here—our algorithm will return states which satisfy a fraction $1/2+AF$ with high probability. If we choose $A$ to be too large compared to the target value established by expressions like Eq. (IV.3.2), we risk failing to well-approximate the problem; conversely, choosing $A$ below it will reduce the returned approximation ratio to $A$ and thus perform suboptimally. We do caution that unlike our predictions for MSFO, this analysis cannot predict the degree of polynomial scaling necessary (and does not formally rule out other sub-exponential but super-polynomial scaling forms). That said, based on the reasoning in section II.5, and numerical testing out to the largest accessible system sizes, it is likely runtimes that scale linearly with $N$ are sufficient. It strikes us as unlikely that superpolynomial decay will survive once Eq. IV.3.2 predicts a positive exponent, in analogy to how for exactly $w=2/3$ the MSFO minimum gap may scale as a stretched exponential, but is firmly in the polynomial scaling regime for $w<2/3$. And, we emphasize again, this prediction assumes a random, potentially dense hypergraph but is fundamentally independent of the fraction satisfied in $E_{\mathrm{GS}}$ itself and so applies to the planted partial solution instances we use for numerical benchmarking below.
IV.3.3 Further Comments and Caveats
@@
We expect that this analysis underestimates the choice of $A$ that will return states with $E\leq AE_{\mathrm{GS}}$ with constant probability. This is because our counting here only counts states very close to the fold, when in reality the probability of tunneling into states a small extensive fraction larger than $AE_{\mathrm{GS}}$ is still going to be appreciable due to the continued exponential growth of the number of targets, even if the per-state tunneling rate does tend to decrease with increasing $E$ due to the interference effect mentioned earlier, in which perturbative corrections that mix with states of lower energy have opposite sign. In addition to this consideration, because the target ground states and low lying excitations of the folded Hamiltonian in $\mathcal{T}$ very roughly form a hyperspherical shell, any individual target state will have other states in $\mathcal{T}$ that are relatively close to it in Hamming distance. Consequently we expect these states to have a band dispersion, centered around the mean energy given by the corrections in Eq. (IV.3.1). Since we are already assuming off-resonant tunneling, i.e. a per-state tunneling rate $\propto\Omega_{0}^{2}$, and so summing over squared matrix elements, if we consider states near the band center where the density is highest this will alter the average tunneling rate by at most a prefactor. However, there are good reasons to suspect that tunneling into extremal states near the bottom of the band can be substantially enhanced, enough to increase the optimal value of $A$. This calculation is difficult to do quantitatively so we do not attempt it here; we instead see the approximation of considering only states near the band center of $\mathcal{T}$ as another choice that likely underestimates the achievable approximation ratio. We discuss this point in more detail in Appendix C.
@@
For quadratic folding AQC as in Eq. (10), if we choose $A=1$ the gap is efficiently computable using the methods in Jörg et al. (2010) and decays as $\Omega_{0}\sim 2^{-0.104N}$. Given that, like linear folding, the cost per flip curve is $A$-independent, if we assume that the tunneling rate per state for $A<1$ is basically equal to this, then the total decay exponent vanishes if $x_{A}\simeq 0.033$ and $A\simeq 0.8$ using the arguments of the previous few paragraphs. We do not think that can be simply assumed as easily as with tunneling between semiclassical minima and a linearly folded problem Hamiltonian, in the DPP, and more theoretical work is needed here to analytically determine the optimal choice of $A$. Interestingly however, our simulation data in Sec. VI roughly supports this conclusion, with a worst case polynomial time approximation ratio of 0.7 found in our simulations. These simulations show that this method performs similarly to, or slightly worse than, the 3XORSAT-TMA algorithm, which is better able to outperform the approximation guarantee of $\sim 0.6$ derived here. [^10] For smaller systems we also tested linear folding AQC and quadratic folding TMA. In very preliminary studies we found that quadratic AQC modestly outperformed linear AQC, and linear TMA more significantly outperformed quadratic TMA. So we chose not to pursue those methods for larger simulations and do not present those results here, but they may be viable or even superior for other problem classes.
Finally, all of the results we obtain here could be further improved by employing a spectral warping transformation to both $H_{P}$ and $H_{L}$, which would reduce the effective degree $p$ toward or even below $2$ (where there is no first order paramagnet-spin glass transition), albeit with the possibility of worsening TFC. In this limit assuming TFC can be ignored the achievable $A$ could become arbitrarily close to 1, as in MSFO. We did not explore this addition in this work, as our goal here was in part to show that folding alone is sufficient for an exponential separation in approximating random instances drawn from $\mathcal{P}\left(N_{C},N,\epsilon\right)$. Coupled with the fact that our simplifying approximations are all likely to underestimate rather than overestimate performance (particularly the band structure, discussed in Appendix C), that a large constant $A$ is achievable without reaching for additional algorithmic tools like warping strikes us as further supporting evidence toward the truth of our core conjecture. [^11] We also expect that the average tunneling rate–and thus, achievable approximation ratio–can likely be further increased by using other, potentially many-frequency, AC methods such as RFQA Kapit (2021b); Kapit and Oganesyan (2021); Tang and Kapit (2021); Mossi et al. (2023); Grattan et al. (2023). For simplicity, we do not incorporate these methods in this work, but they could be a novel way to further improve the performance of this algorithm and are worth exploring in future research.
@@
In summary, through a relatively novel resummed extensive order perturbation theory, we have shown that random hypergraph MAX-3-XORSAT instances, including extremal ones with planted partial solutions, are efficiently approximable to a fairly large constant fraction through spectrally filtered quantum optimization algorithms. We do not expect this to be the case for TAQC (or QAOA), for the reasons discussed earlier in section II and supported by the numerical evidence we present below. We similarly do not expect such guarantees to be possible for directly finding global optima (particularly as $d_{C}$ becomes small or as $\epsilon$ approaches 1/2), for the reasons set forth in the introduction, though in practice MSFO can often find $G$ in polynomial time when $d_{C}$ is large and $\epsilon$ is small. Evidently, one cannot so easily summit a mountain in hyperspace, but one can reach the rim of a crater. We now present a series of numerical simulations to further support these claims.
@@ V.1 Setup and summary of results
To confirm our predictions–or at least, verify that any serious issues with our calculations and interpretation of the problem are subtle and not apparent at system sizes within reach of present or near-future classical simulations–we performed a series of numerical simulations of various classical and quantum algorithms applied to our PPSPs. For all quantum simulation tasks we used the Qulacs package Suzuki et al. (2021). For smaller systems and algorithm prototyping we ran our simulations on local workstations and cluster nodes; this includes all the spectral folding TMA simulations. For all TAQC and spectral folding AQC simulations presented, we used the Fujitsu Quantum Simulator, a classical HPC system. This allowed us to probe larger system sizes while still averaging over enough instances to have reliable statistics for the problem class. Unless otherwise stated, each datapoint represents the average over 960 or 1000 (for the TAQC and spectral folding AQC results) random problem instances constructed with the prescriptions outlined in section II.4. In all cases in this work we used an unsatisfied fraction $\epsilon=0.1$ for our partial planted solutions; we expect similar phenomenological behavior for other constant $\epsilon$ values, though the thresholds we measure would vary with $\epsilon$ for classical optimization algorithms and TAQC (but not folded optimization, below its predicted performance floor).
@@
To determined the polynomial time hardness threshold, we applied a very simple rule where we fitted $P_{q}\left(N\right)$ to a simple exponential function, assuming any observed decay corresponded asymptotically to exponential decay, and any positive exponents, e.g. exponential growth, represent small-$N$ growth toward some constant saturation value. For the greedy algorithm with $N_{C}/N=1.5\sqrt{N}$ we fitted decay to an exponential in $\sqrt{N}$ as discussed below. The polynomial time approximation hardness threshold for a given algorithm, unsatisfied fraction $\epsilon$ and $N_{C}/N$ scaling choice is defined to be $q_{a}$, the largest values of $q$ for which we do not observe exponential decay. For $N_{C}/N=3\sqrt{N}/2$ this threshold is decaying with system size, and the asterisk next to the result for TAQC is to highlight the fact that we only ran these simulations out to $N=28$ so are likely not capturing the asymptotic threshold.
This notation highlights a difference from $q_{m}$ in our core conjecture; $q_{a}$ corresponds to the best achievable approximation ratio for a given algorithm and problem parametrization, where $q_{m}$ is the global minimum value of $q_{a}$ across all spectrally filtered optimization variations and problem parametrizations (again restricted to random instances drawn from $\mathcal{P}\left(N_{C},N,\epsilon\right)$). We also want to emphasize that $q_{a}$ is a measure of the performance of an algorithm, whereas the approximation target $A$ in folded optimization is a control parameter, and $q_{a}$ may well fall short of it. As discussed elsewhere, the number of cost function calls is $O\left(N\right)$ for all approaches studied , except MSFO. Cases for which a range is quoted are where we felt there was some ambiguity to the fitting, and all values are the result of extrapolating fits to numerical simulations and are naturally somewhat approximate.
$N_{C}/N$ Classical TAQC AQC/0.75 AQC/0.85 TMA-3/0.75 TMA-3/0.85 TMA-L/0.75 MSFO
2 0.75-0.8 0.75 0.75 0.75 0.75 0.75 0.7 1*
4 0.55 0.55 0.7 0.7 0.75 0.8 0.75 1*
6 0.45 0.45 0.75 0.7 0.8 0.8 0.75 1*
3 $\sqrt{N}/2$ decaying 0.25/decaying* 0.75 0.75 0.8 0.8 0.75-0.8 1*
@@
Table 1: Approximation hardness thresholds for the classical greedy search, high-depth TAQC and spectral folding variations. This table lists $q_{a}$, the largest value of the approximation ratio $q$ before exponential decay is reported, drawn from the numerical experiments in figures 7 through 11 and 18. Spectral folding results are labeled as protocol/$A$ (where $A$ is the approximation target); AQC is quadratic spectral folding in the AQC formulation, TMA-3 is trial minimum annealing with a linear folded Hamiltonian and 3-XORSAT lowering Hamiltonian, and TMA-L is the same with local $Z$ biases for the lowering Hamiltonian. MSFO is the multi-stage filtered optimization algorithm of section III.3; using a runtime that scales approximately as $N^{3/2}$ (in contrast to the linear runtimes of all other methods) it is capable of finding the ground state directly with constant probability, at least for random instances in the parameter ranges simulated. These results–where the threshold decays for traditional methods but not folded optimization–support the predictions in section IV that random hypergraph problems are efficiently approximable through spectrally filtered quantum optimization.
V.2 Performance of quasi-greedy classical algorithms